find all zeros of the polynomial f(x)=x^3+9x^2-36x-324 and determine the multiplicity of each?
Answer by ewatrrr(24785)

 
Hi,
f(x)=x^3+9x^2-36x-324
f(x)=x^2(x+9)-36(x+9) ***
f(x) = (x^2-36)(x+9)
If f(x) = (x-6)(x+6)(x+9)= 0
x is -9, -6, 6